OpenStudy (anonymous):

3. Solve the equation and write the solution set. |2x – 6| – 4x = 8x + 4

OpenStudy (anonymous):

this is the only question i cant seem to find the right answer for! can someone please help?

OpenStudy (the_fizicx99):

Do you have options to chose from?

OpenStudy (anonymous):

no, thats all it gave me.

OpenStudy (the_fizicx99):

kk, sec

OpenStudy (the_fizicx99):

I got -10x/10 = 10/-10

OpenStudy (anonymous):

x=-1

OpenStudy (the_fizicx99):

Yup :D

OpenStudy (anonymous):

okay thanks! :) and how did you get that?

OpenStudy (the_fizicx99):

l2x - 6l -4x = 8x + 4 + 4x + 4x 2x - 6 = 12x + 4 + 6 + 6 2x = 12x + 10 -12x - 12x -10x = 10 /-10 /-10 = -1

OpenStudy (anonymous):

well done fizzy...and sara i hope ths guy let u to understand thorouly

OpenStudy (the_fizicx99):

Fizzy? =.=

OpenStudy (the_fizicx99):

@sarahann117 did you understand?

OpenStudy (anonymous):

fizzi which is the best free software for mathmetics,i e Calculus,fouriour n Laplace Transforms???

OpenStudy (the_fizicx99):

I'm a freshman lol I can't take calculus yet.. sorry :/

OpenStudy (anonymous):

ok fizzy in which level r u studying?

OpenStudy (the_fizicx99):

What do you mean?

OpenStudy (anonymous):

i mean in which class r u studying?

OpenStudy (the_fizicx99):

Algebra I. though I practice more advance stuff when I'm not doing HW. I'm kinda learning physics too :)

OpenStudy (the_fizicx99):

This helped me: https://www.mathway.com GL :>

OpenStudy (anonymous):

ok but i am studying advance Engineering Mathmetics ...only applied in Electrical Engineering

OpenStudy (the_fizicx99):

Have you looked at mathway.com?

OpenStudy (anonymous):

yes it is good website but u can also try integrals.wolfram.com/‎ and Quickmath.com as well

OpenStudy (anonymous):

Yes I've done a few other problems like this one but i forgot to carry over, thanks though!(:

OpenStudy (the_fizicx99):

Yea, saw them on google :D @sarahann117 Yw :)

OpenStudy (anonymous):

ok dear thanks u so much,,,if u are using facebook too here is my id skywalker763@yahoo.com u can add me their

OpenStudy (the_fizicx99):

xD sure .. I forgot my fb password and possible email ._.

OpenStudy (anonymous):

ok i hope u ll add whenever u ll make ur new account on fb

OpenStudy (the_fizicx99):

kk cyaa :p

OpenStudy (math&ing001):

@sarahann17 This kind of equations requires taking in consideration intervals of x for the absolute value of the expression |2x-6| . Case 1: when 2x-6>=0 ie. x>=3 The answer is like @tHe_FiZiCx99 and @hasnain986 said : x=-1 But -1<3, meaning the answer doesn't belong in the interval. Therefore, there's no solution on this interval. Case 2: when 2x-6<=0 ie x<=3 |2x-6| - 4x = 8x + 4 becomes -(2x-6) - 4x = 8x + 4 The answer to that is x=1/7 And 1/7<=3, meaning the answer belongs in the interval. So the solution set: S={1/7}

OpenStudy (math&ing001):

@sarahann117 *

OpenStudy (the_fizicx99):

I also got 1/7 when I tried to find the interval. forgot to include it though Ty @math&ing001